OpenStudy (anonymous):

What is the lateral area of a rectangular prism if the base edges are 8 feet and 6 feet and the height is 10 feet? A. 48 ft2 B. 280 ft2 C. 480 ft2 D. 240 ft2

OpenStudy (ganpat):

Answer is 280.. Lateral area for rect = p * h where p = perimeter = 2(6 +8) and h = height = 10

OpenStudy (anonymous):

Ok thanks!

OpenStudy (ganpat):

anytime..